OpenStudy (anonymous):

i think i am being overwhelmed by a large number. i am looking to first solve by clearing decimals, then i want an answer for x. .06(10,000) + .08x = .072 (10,000 + x)

OpenStudy (anonymous):

NO worries .06(10,000) + .08x = .072 (10,000 + x) make life easier by multiplying 1000 throughout the whole equation resulting in , 600 000+80x = 720 000 +72x 80x-72x = 720 000 - 600 000 8x = 120 000 x = 15 000

OpenStudy (anonymous):

ok, so i think i forgot to distribute the 072 times x in the second half! thank you!!!! sigh... two hours right down the drain that i can't get back!! your a life saver!!!

OpenStudy (anonymous):

how would i convert 7/32 to a decimal?